Pixel:

If your medical bill is $624 and you have 80% medical coverage, $150 deductible, how much will you have to pay?

Pixel:

624 x .8 = 499.20

Pixel:

624 - 499.20 =124.80 124.80 + 150 = 274.80

Pixel:

i think i did this wrong

TheSmartOne:

I don't have that much experience with insurance so I might be completely wrong But wouldn't you pay the deductible first and then the 80% coverage kicks in?

Pixel:

624 - 150 = 474 474 x .8 = 379.2

TheSmartOne:

Because that's how deductibles work. Like what if it was a $170 cost. If you do your method, it would be like weird Also instead of multiplying by 0.8 and then subtracting, you can just multiply by 0.2 Will save you an extra calculation ;)

TheSmartOne:

So yes, you paid $150 And you're left with 474, out of which the insurance will pay 80% of it which is 379.2 You could do it two ways, which is what I was trying to explain earlier: how you could save one step 474 * 0.2 Or 474* 0.8 = some 'x' And then 474 - x You'll see that you get the same answer for both

ASpopa182005:

You pay $150 + 20% of amount of bill above $150 $150 + 20% ($624−$150) = $150 + 0.2 ($474) = $244.80 The answer is $244.80

Pixel:

so its 379.2

TheSmartOne:

Ye what @ASpopa182005 did, also welcome to QC! We value your willingness to help others. I'd just like to redirect you to our rules https://questioncove.com/rules We take a different approach, where we try to teach instead of giving the answer

TheSmartOne:

379.2 is just 80% after the deductible that the insurance pays Not what you end up paying in all for the medical bill